Calculate the standard entropy, Asron, of the following reaction at 25.0 oC using the data in this table. The standard enthalpy of the reaction, rXn is -633.1 kJ mol AH 3C,H, Number J. K 1. mol Then, calculate the standard Gibbs free energy of the reaction, AGoron. Number kJ mol AG. Finally, determine which direction the reaction is spontaneous as written at 25.0 eC and standard pressure. O forward O reverse O both O neither
Answer #1

For reaction we need:

SRxn = Sproduct - Sreactants

Srxn = Benzen e- 3*Ethylene = 173.4 -3*219.4 = -484.8J/molK

b)

Grxn:

dG = dH - T*dS

dG = -633.1*1000 - 298*(-484.8) = -488629.6 J/mol = -488.63 kJ/mol

c)

this is negative in dG change, so the process must be SPONTENEOUS